IODP Proceedings    Volume contents     Search



Our objective at Site U1383 was to install a multilevel CORK to perform long-term coupled microbiological, biogeochemical, and hydrological experiments in deeper portions of the oceanic crust. The operational tasks conducted at Site U1383 are given in Table T1. A summary of each of the three holes occupied at Site U1383 is presented in Table T2. Our first step was to use an 18½ inch tricone drill bit to perform a jet-in test to determine the length of 20 inch casing to install with the reentry cone. While we assembled this bottom-hole assembly (BHA) and lowered to the seafloor, we retrieved the seafloor beacon used for Hole 395A and transited ~5.8 km to Site U1383 in dynamic positioning mode. After arriving at Site U1383, we dropped a positioning beacon at 1812 h on 12 October 2011 (all times are local ship time, Universal Time Coordinated [UTC] – 3 h). The jet-in test in Hole U1383A began at 2240 h on 12 October and reached 36 mbsf at 0200 h on 13 October. The penetration rate during the last 2 m was very slow. The 18½ inch bit was back on board at 0930 h, ending Hole U1383A.

On the basis of the jet-in test, we deployed 34.84 m of 20 inch casing for Hole U1383B. The casing and reentry cone were run to the bottom, but our initial attempt jetted in the casing to only 29 mbsf. After trying to advance the casing for over 3 h with no progress, we pulled the string out of the seafloor and moved 50 m to the northeast. Our second attempt to start Hole U1383B began at 0810 h on 14 October. This time, we were able to fully land the reentry system and casing at 1015 h. We released the running tool at 1035 h and retrieved the drill string, with the bit arriving on the rig floor at 1915 h.

Our next job was to drill an 18½ inch hole into uppermost basement for the 16 inch casing. After both drill collar pup joints and the casing running tool were laid out, the 18½ inch bit was removed, the nozzles were changed, and the BHA was made up and run back to the bottom. During the trip into the hole we had to pause to slip and cut 115 ft of drill line and to start lowering the camera system for reentry. We reentered Hole U1383B at 0518 h on 15 October, began drilling at 0704 h, and encountered basement at 0723 h at 53 mbsf. Drilling continued to 68 mbsf at a penetration rate of ~2 m/h. We drilled to this depth to allow ~9 m of rathole for the 16 inch casing. Once 68 mbsf was reached, the hole was reamed and conditioned and the bit was retrieved.

After the bit was back on the rig floor at 0450 h on 16 October, the casing running tool was picked up and made up to the BHA and set back in the derrick. A total of 58.82 m of 16 inch casing, including the casing hanger, was assembled. The running tool was made up to the casing, and the casing was then run to the bottom, pausing only to deploy the camera system. Hole U1383B was reentered at 1524 h after ~20 min of maneuvering. The casing was then lowered into the open hole until the casing shoe contacted the sediment/​basement interface. The casing shoe would not pass the interface after 1 h of attempts, even with rotation. The ship was offset, and repeated attempts were made until the shoe passed the basement contact. Unfortunately, the casing was unable to advance enough to successfully land the casing hanger in the reentry cone. The final position of the casing shoe was ~3 m short of landing. The 16 inch casing was retrieved at 1945 h on 16 October, arriving back on the rig floor at 0315 h on 17 October. The top two joints of casing were removed, and the length of the casing shoe joint was shortened by 5 m. A new casing shoe was welded onto the shoe joint, and at 0800 h the casing was lowered back into the hole. The running tool was reattached, and the casing was once again run back to the bottom. The camera system was installed 70 stands into the trip to the bottom. At 1537 h, Hole U1383B was reentered with the 16 inch casing for the second time. The casing was successfully landed at 1630 h and cemented into place with 20 bbl of cement with lost-circulation material (LCM; Cello-Flake). The casing running tool was released at 1745 h, and the drill string was circulated out to remove any excess cement from the string. At 1845 h on 17 October, the drill string was pulled back to the surface, stopping only so that the camera system could be removed. The running tool cleared the rotary table at 0130 h on 18 October and then was detorqued and laid out.

The next stage of operations was to drill ~100 m of basement to install 10¾ inch casing. A new 14¾ inch bit was made up to the BHA, and the drill string was tripped to the bottom. At 90 stands into the trip, the camera system was installed and run to the bottom behind the bit. The trip was suspended just above the seafloor to slip and cut 115 ft of drilling line from the drawworks winch. Hole U1383B was reentered at 1210 h on 18 October after 15 min of maneuvering. The bit was tripped to near the bottom, the top drive was picked up, and cement was tagged at 49 mbsf (5 m above the casing shoe). After 45 min of drilling, the bit reached the bottom of the cement at 57 mbsf (3 m below the casing shoe). From 1500 h on 18 October through 0515 h on 19 October, drilling proceeded to 89.8 mbsf. At this depth, torque increased, revolutions per minute became erratic, and penetration rates dropped to zero. A wiper trip was performed from 89.8 mbsf to the 16 inch casing shoe at 54 mbsf and back to the bottom. Attempts to resume drilling were unsuccessful. At 1000 h, we decided to pull the bit and inspect it at the surface. The top drive was set back, and the drill string was pulled from the hole. The bit cleared the rotary table at 1750 h on 19 October. Inspection of the bit revealed that two of the three rotary cones and both shanks, including bearings and nozzles, had broken off the body of the bit. Only one cone remained on the bit, and it too had all the inserts broken off. After discussions with all of the operations staff on board, we decided that we would have a greater chance of achieving the expedition objectives by starting a new hole rather than trying to salvage Hole U1383B. The final configuration of reentry cone, casing, and hole depth is shown in Figure F1. Subsequently, we deployed a ROV platform on top of the reentry cone (see below).

At 1830 h, construction began on a reentry cone for Hole U1383C. Because of several factors, we decided to start the new reentry system with 16 inch casing jetted in to 34.58 mbsf at the same location as the original jet-in test in Hole U1383A. Assembly of the Hole U1383C reentry cone was completed at 0630 h on 20 October. The reentry cone was moved to the moonpool and positioned on the moonpool doors, and then 35 m of 16 inch casing was picked up, landed, and latched into the reentry cone. The running tool was then released, and the stand was set back in the derrick. The stand with the 14¾ inch drill bit was then picked up, followed by another 5 m of 8¼ inch drill collar pup joints to space the bit out to the 16 inch casing shoe. The drill collar stand with the casing running tool was picked up and made up to the BHA and lowered and latched into the 16 inch casing hanger. The remainder of the BHA was assembled, and the reentry system and casing were lowered to just above the seafloor, stopping every 30 stands to fill the drill string. The camera system was installed 90 stands in and run to just above the reentry system. Hole U1383C was spudded at 1900 h on 20 October. The reentry system was jetted in ~35 m, and the casing running tool was released at 2130 h. The camera system and drill string were then pulled back to the surface, and the 14¾ inch bit cleared the rotary table at 0545 h on 21 October. After the nozzles on the bit were changed, the drilling BHA was assembled and run back to the bottom. The camera system was installed at Stand 70 and run to the bottom, following the bit to the seafloor. At 1130 h, the vessel began maneuvering, and Hole U1383C was reentered at 1444 h. After the top drive was picked up, drilling began at 1635 h on 21 October. Contact was made with basement at 38.3 mbsf, just 3.6 m below the 16 inch casing shoe, after only 10 min of drilling. Basement drilling continued with slow penetration rates (0.5–2.5 m/h), reaching a total depth of 69.5 mbsf at 1840 h on 22 October.

We assembled 60.41 m of 10¾ inch casing with a swellable packer just below the casing hanger and lowered it to just above the seafloor, pausing about halfway to deploy the camera system. After ~30 min of maneuvering, we reentered Hole U1383C at 1626 h on 23 October. The casing string was landed at 1720 h, and we confirmed it was latched in with a 20,000 lb overpull. The casing was secured in place by pumping 25 bbl of cement with LCM (Cello-Flake). The casing running tool was released at 1940 h, the drill string was flushed clear of any remaining cement, and we started pulling out of the hole. With the casing running tool at 2336 meters below rig floor (mbrf), the trip was halted so we could slip and cut 115 ft of drill line. We also spent 2 h replacing a spool valve and repairing a hydraulic line on the 5 inch pipe racker. We then continued retrieving the casing running tool, which was back on board at 0825 h on 24 October.

We assembled a RCB bit and BHA, verified the core barrel space out, and lowered it to the seafloor. About halfway down, we paused to deploy the camera system. This reentry took us only ~10 min of maneuvering. After we retrieved the camera system, installed the top drive, and spaced out for drilling, we tagged the top of the cement at 43.8 mbsf (~16.6 m above the casing shoe). Once we drilled out the cement and cleaned out the rathole, we recovered the center bit. We dropped an RCB core barrel and started coring at 0000 h on 25 October. A summary of the cores recovered in Hole U1383C is presented in Table T2. We cut Cores 336-U1383C-2R through 18R from 69.5 to 211.6 mbsf. After Core 15R had been cut, we performed a wiper trip back up to the casing shoe and encountered 20 m of fill while getting back to the bottom of the hole. Hole cleaning remained a priority, with mud sweeps being performed on average twice during each cored interval. A total of 142.1 m was cored, and 28.55 m was recovered (20%). Just as Core 18R began to be cut (with 45.8 h on the bit), the vessel began experiencing high heave. This heave made it almost impossible to keep sufficient weight on the bit to keep it on the bottom, so we circulated cuttings out of the hole with a final mud sweep and pulled the bit out of the hole. The RCB bit was back on board at 1720 h on 28 October. It was fortunate that we did not continue coring with this bit because it had experienced bearing failures on all four roller cones.

We conducted routine rig maintenance while we assembled a new C-7 RCB bit and inspected the float valve and support bearing assemblies. This maintenance identified a faulty air cylinder for the locking pin on the motion compensator that had to be repaired, which resulted in 9.25 h of rig downtime. During the repair, we assembled the previous BHA and added three more drill collars. We installed a center bit so that the float valve would remain open during the trip, allowing the drill string to fill with seawater. At 0630 on 29 October, we started tripping the BHA to the seafloor. The reentry cone was visible as soon as the camera system reached the seafloor, and the bit was almost directly over the cone. We reentered Hole U1383C in several minutes at 1231 h on 29 October. We retrieved the camera system and the center bit and lowered the bit to the bottom of the hole to check for tight spots and to circulate out any fill at the bottom before changing bits.

We assembled a second RCB bit to continue coring in Hole U1383C. After the bit was lowered to the bottom of the hole, it was circulated clean. An RCB core barrel was dropped, and coring resumed with Core 336-U1383C-19R. We decided to core with drilling knobbies because of high vessel motion and operating in deep water. RCB coring continued through Core 336-U1383C-23R, at which point we made a short wiper trip to both clean a short section of hole and to replace three drilling knobbies with a stand of 5½ inch drill pipe. Once the bit was back on the bottom, we circulated a mud sweep and resumed coring. Two 20 bbl sweeps were pumped on every core to keep the hole clean. We decided to stop RCB coring at 4756.7 mbrf (331.5 mbsf) because of accumulating bit hours, minimal time left in the expedition, and because a sufficient depth was reached to achieve the CORK objectives. Thirty-one cores recovered 50.31 m of rocks from a 262.0 m interval. Overall recovery was 19%, varying from 8% to a maximum of 43%.

We conducted three wiper trips to prepare the hole for downhole logging, packer experiments, and CORK installation. At 1400 h on 1 November, the first wiper trip began, and hole cleaning continued until the following day. Three wiper trips were made from total depth to the 10¾ inch casing shoe and back. Tight spots were recorded at 4692, 4703, 4708, 4716, 4724, 4743, 4751, and 4752 mbrf (266.8, 277.8, 282.8, 290.8, 298.8, 317.8, 325.8, and 326.8 mbsf). These spots were reamed and rechecked during the first wiper trip. At the end of the first wiper trip, 4 m of fill was found on the bottom and cleaned out. We circulated a 70 bbl sweep of high-viscosity mud, and then the hole was displaced with salt water. The subsequent two wiper trips did not detect tight spots or fill on the bottom. At 0545 h on 2 November, we started to pull the string out of Hole U1383C, and it cleared the seafloor at 0625 h. The RCB bit was back on the rig floor at 1315 h.

Before we started assembling the logging/​packer BHA, we picked up and assembled six 6¾ inch perforated drill collars into a stand. The exterior of the stand was then completely coated with an epoxy paint and set back in the derrick to cure. This stand is the lowermost portion of the CORK completion string.

The logging and drill string packer BHA was then assembled and lowered to the seafloor. After ~3 min of maneuvering, we reentered Hole U1383C at 2342 h on 2 November. The logging bit was positioned ~5 m from the base of the 10¾ inch casing shoe.

The logging equipment was rigged up to run the adapted microbiology combination II (AMC II) tool string, which included the DEBI-t deep UV–induced fluorescence, density/​caliper (Hostile Environment Litho-Density Sonde [HLDS]), and spectral gamma ray (Hostile Environment Natural Gamma Ray Sonde [HNGS]) tools. During assembly, the resistivity tool failed and had to be removed from the string. The tool string was run into the hole, and two full passes were conducted, tagging bottom and confirming drill pipe depth and lack of fill in the hole. After this first tool string was back on board, the FMS-sonic tool string was rigged up. This tool string included the spectral natural gamma ray, Dipole Shear Sonic Imager (DSI), and FMS tools. This second string also completed two successful passes; it was pulled out of the hole and rigged down at 2315 h on 3 November.

Following logging, we picked up the top drive and prepared for the first of three scheduled packer tests. We were unable to set the packer in the casing, and the test at this level was abandoned. The packer was then moved to 4566 mbrf (140.8 mbsf) in open hole. The packer was inflated, and flow tests were completed as scheduled. The packer was then moved to the final position but was unable to hold pressure. The packer experiment was concluded, the top drive was set back, and the drill string was pulled out of the hole to just above the seafloor.

Before beginning CORK installation in Hole U1383C, we decided to deploy an ROV platform in Hole U1383B to facilitate a future borehole observatory installation by ROV. We moved to Hole U1383B, deployed the camera system, and reentered the hole at 1350 h on 4 November. We retrieved the camera system, assembled a modified ROV platform around the drill pipe, and let it free fall down onto the reentry cone at the seafloor. The platform had been modified to self-center on the reentry cone because there was no CORK wellhead to help it center as designed. The camera was deployed, and the platform was observed to be sitting in the cone—although perhaps just slightly off-center. The bit was pulled out of the hole at 1818 h. After pulling the bit well above the seafloor, we paused operations to slip and cut the drill line. At 2015 h, the trip out of the hole resumed, and the logging bit arrived back on the rig floor at 0300 h on 5 November. We observed that the drill string packer element had experienced a blowout. After we had the drill collars stored in the derrick, we began assembling the Hole U1383C CORK.

A schematic of the installed observatory is shown in Figures F2 and F3. The individual CORK pieces assembled are listed in Table T3. The observatory assembly started with lowering the preassembled 6¾ inch perforated drill collar stand below the rig floor. We picked up, assembled, and painted with epoxy four additional 15 ft long perforated drill collars, which we added to the top of the first stand to provide additional weight to the bottom of the CORK to hopefully ease the string past a ledge encountered during hole cleaning. We then started assembling the various pieces of the CORK, including (1) external umbilicals terminated at screens, (2) coated and perforated steel and slotted fiberglass tubing, (3) landing seats, (4) packers, and (5) a variety of required crossovers. Five external umbilicals allow access to these three zones for microbiological, geochemical, and pressure sampling. The CORK tubing (coated steel and fiberglass) extends to 247.6 mbsf and includes perforated and slotted intervals (76.3–129.4, 145.9–181.1, and 203.7–246.6 mbsf) that provide access to the three isolated intervals but leave the bottom portion of the hole open for future logging and access (247.6–331.5 mbsf). An internal OsmoSampler string extends the full length of the CORK and includes seals isolating the three zones and microbiological, geochemical, pressure, and temperature experiments.

At just after 1700 h on 5 November, we finished assembling the 247.6 m long CORK observatory and picked up the CORK head and attached it to the casing string. The last umbilical connections were made to the bottom of the CORK head. The CORK was lowered to the moonpool and landed on the moonpool doors. The CORK running tool was removed, and the OsmoSampler instrument string was assembled and lowered inside the CORK and casing. The CORK running tool was reinstalled, and the packer inflation line was connected. The CORK was then lowered below the keel to flood the umbilicals to remove air from the pressure lines. The CORK was then raised to the moonpool, and the seafloor fast-flow OsmoSampler was installed. All valves except the geochemistry bay upper Zone 2 valve (which was connected to the seafloor fast-flow OsmoSampler) were closed and secured. The CORK was lowered to ~100 meters below sea level (mbsl), and the camera frame was test fit over the CORK head. At 0000 h on 6 November 2011, after ~21 h of assembly, we started lowering the CORK assembly to the seafloor. During the trip in the hole, the drill pipe was filled with seawater every six stands, which lengthened the trip to a total of 8.25 h. At 0915 h, the vessel began maneuvering to reenter Hole U1383C. The hole was reentered at 0941 h. After reentry, the CORK was slowly lowered into the hole while drill string weight was carefully monitored and the procedure was observed with the camera. Approximately 35 m above the landing point, the top drive was picked up, and the CORK was fully landed at 1127 h on 6 November. Afterward, the mud pumps were brought online to pump up the packers. We applied pressure to the packers for 30 min but were unable to get them to hold a sustained pressure. It was evident there was a small leak somewhere in the system between the standpipe on the rig floor and the bottom packer. We were unable to detect where the leak might be, and there are many possibilities or combinations depending on where the leak is—from all packers pressurized and sealing to only the top swellable packer sealing. After the camera was recovered, the ROV platform was assembled, hung from the camera frame, and run to the bottom. The platform was released at 1630 h without incident. The camera was pulled back to the surface, the slings and releasing tool were removed, and the camera was lowered back to the bottom to monitor the release of the running tool from the CORK head. The running tool was released at 2015 h on 6 November, ending the Hole U1383C CORK deployment. When the CORK running tool was released, the CORK wellhead also appeared to rotate. Several possibilities could account for this action, but it is likely the result of tightening the upper casing in excess of the manual tightening on the rig floor. A similar occurrence was noted with the Juan de Fuca CORKs. The camera and drill string were retrieved, with the CORK running tool arriving back on the rig floor at 0315 h on 7 November, ending Hole U1383C.